A golfer wants to drive a golf ball a distance of 310yards (283m).  If the 4-wood launches the ball at 21° above the horizontal, what must the initial speed of the ball be in order to achieve the required distance?


1
Expert's answer
2021-05-12T10:08:23-0400

Answer

Vertical velocity can be obtain as

"(vsin21°)t-\frac{1}{2}(9.8)t=0\\
vsin21°-4.9t=0\quad... Eq. 1"


Horizontal velocity can be obtain as


"vcos21°t=283\\t=\frac{303.1}{v}"


By eq. 1

"0.36v-\frac{1485.2}{v}=0\\v=64.23m/s"